OpenStudy (seba_beto):

If you bet on “red” at roulette, you have chance 18/38 of winning. Suppose you make a sequence of independent bets on “red” at roulette, with the decision that you will stop playing once you’ve won 5 times. What is the chance that after 15 bets you are still playing?

OpenStudy (perl):

ok ill try again but

OpenStudy (perl):

do you have the answer to check ?

OpenStudy (seba_beto):

I have the asw. my problem is in understanding lol

OpenStudy (perl):

i just need to know it is right before explaining it

OpenStudy (seba_beto):

yeah me too!

OpenStudy (seba_beto):

it is

OpenStudy (perl):

\[P(x <5)=\sum_{k=0}^{4}\left(\begin{matrix}15 \\ k\end{matrix}\right)\left( \frac{ 18 }{ 38 } \right)^k \left(\frac{ 20 }{38 }\right)^{15-k}\]

OpenStudy (perl):

makes sense?

OpenStudy (seba_beto):

sort of...

OpenStudy (seba_beto):

I've got it!!! it's just a binomial proba. thx